Suppose a consumer has income of $120 per period and faces prices pX = 2 and pZ = 3. Her goal is to maximize her utility, described by the function U = 10X0.5Z0.5. Calculate the utility maximizing bundle (X* , Z* )
Answer #1

] C-10x² z ² + x[120-2x-32 6 Px=2 P2=3 I = 120 ah - Llox 1 1 2 -2 = 1) 24 = 1.10x625 - 34 = 0 -- 26 - 120 = 2X +32 . ☺ . 24 D